# Optimization of Module Transferability in Single Image Super-Resolution: Universality Assessment and Cycle Residual Blocks

### v1

@@ 4.2 Universality Assessment for Diverse Modules
Table 2 demonstrates the core determinants of universality. Although RCAB [11] possesses larger $n$ and $l$ compared to RSTB [13], the elevated $k$ in RSTB deteriorates its transferability. This finding aligns with empirical evidence: RSTB integrates three nested submodules—MSA, MLP, and STL—whereas RCAB employs only a CALayer. Hence, effective migration of RSTB requires the concurrent adaptation of multiple components. Conversely, despite $n_{\mathrm{CRB}}\gg n_{\mathrm{DCRB}}$, DCRB remains more difficult to transfer owing to the inclusion of depth-wise convolution, even though at a macroscopic level $\phi(\text{DCRB})\approx\phi(\text{CRB})$.
@@
Figure 9: Quanlitative comparison of coalmine images for ×$\times 4$ SISR.
Figure 10: Qualitative comparison of remote sensing images for ×$\times 4$ SISR.
@@ 4.3 Computational Overhead Analysis
As illustrated in Fig.7, CRB achieves the highest PSNR, while DCRB delivers comparably strong fidelity with the lowest FLOPs, parameter count, and GPU runtime. Furthermore, CRB outperforms RB in PSNR under equivalent parameters (tested on Potsdam [43]). The convergence profiles in Fig.7 indicate that all universal modules exhibit similar training patterns, with CRB attaining marginally faster convergence within the first 20 epochs, a behavior attributable to the accelerated gradient flow afforded by its additional residual branches (Eq.(7)).
@@ 4.4 SISR for General Scene Images
Using EDSR [6] as the backbone network, we conduct a combined analysis of module universality and performance, as shown in Table 3. Nearly all of the optimized modules feature better generalizability than baseline RB by slightly sacrificing their universality. Notably, RCAB–despite suffering the greatest universality degradation–fails to achieve the best performance. In contrast, CRB attains the largest generalization gains with minimal impact on universality, marking it as a positively universal module that could be integrated into diverse architectures without additional computational overhead. On the one hand, this finding demonstrates that CRB represents an exceptionally ideal module. On the other hand, this finding indicates that universality and generalizability are not isolated. Although prior work has shown that higher generalization could be attained by sacrificing module transferability—via increased module complexity and representational capacity—their precise interplay remains contingent on the specific architecture of plug-and-play modules.
Figure 11: Hardware deployments and the practical applicability evaluation of optimized modules.
Table 4 summarizes CRB and DCRB performance across multiple SR scales. EDSR augmented by CRB consistently achieves the highest fidelity (PSNR/SSIM), whereas DCRB yields only minor performance degradations—-never exceeding a 5% gap and reaching a minimum of 0.23% under ×$\times 4$ SR on B100. Remarkably, DCRB comprises just 1.73% of the parameter count of CRB and RB (Table 2), yet under certain conditions delivers approximately 99% of the baseline performance.
As illustrated in Fig.8, we present the ×$\times 3$ SR results on the B100 dataset [36] to visually compare CRB (Ours1) and DCRB (Ours2). Both variants achieve marked clarity enhancements over the LR, yielding similar qualitative reconstructions. The residual images-—defined as the per-pixel discrepancy between HR and SR images-—reveal that CRB substantially reduces errors in critical details (e.g., roof eaves in 97033.png). In contrast, DCRB incurs increased residual magnitudes compared to the baseline, particularly within detailed texture areas where errors are more pronounced.
@@ 4.5 SISR For Remote Sensing Images
Table 5 reports the remote-sensing SISR results. Models augmented with CRB consistently outperform their baseline counterparts, demonstrating CRB’s superior universality and its positive impact on model generalizability. In particular, “IPG$+$CRB” achieves the best performance. In contrast, DCRB-—owing to its depthwise separable convolution design-—occasionally underperforms CRB.
@@
4.6 SISR For Industrial Scenes: Applications on Low-Light Mine Imagery Reconstruction
Using the BelT dataset [44], we evaluate the optimized models’ performance in extreme industrial settings (e.g., mine environments), thereby validating the effectiveness of CRB and DCRB.
To reconstruct low-light, low-resolution mine imagery, we cascade KinD [45], the illumination-enhancement module, at the model’s input and perform SISR under normal illumination at the output. Results are summarized in Table 6. Nearly all of the models embedded with CRB/DCRB witness improvements over the baseline while IPG augmented by CRB achieves the best performance.
Fig.LABEL:coalmine_1 provides qualitative evidence: In Image 1, the mechanical structures flanking the belt are reconstructed with sharply defined edges. In Image 2, the transported materials, initially imperceptible, become distinctly visible after reconstruction. In Fig.LABEL:coalmine_2, we plot the pixel intensities along the 125th row of SR outputs. All reconstructed channels show a pronounced increase in amplitude, with CRB yielding marginally higher values than DCRB. This confirms that the optimized modules preserve original image features while amplifying high-frequency details, thus improving both signal fidelity and perceptual realism.
4.7 Hardware Implementations and Extra low-light Applications
In addition to software optimizations, we conduct on-device hardware validation of CRB and DCRB modules using Jetson Nano B01 paired with a CSI camera to assess their practical applicability. The hardware setup is illustrated in Fig.LABEL:hardware.
Fig.LABEL:hardware_results provides hardware-based evidence supporting the effectiveness of proposed modules, where the optimized model successfully reconstructs low-light&low-resolution images that KinD alone cannot reconstruct, thereby markedly enhancing the critical details (e.g., object textures).
From the spectral density plots in Fig.LABEL:hardware_results, it is evident that, unlike the LR image whose spectrum is dominated by an intensely bright central region indicating redundant low-frequency information, the SR image recovers substantially more high-frequency details. This is illustrated by the SR spectral density’s overall brightness extending toward the high-frequency regions and similarity to the HR spectrum. Although a slight haze remains in the reconstructed output, it does not compromise the overall visual quality.

### v2

@@ 4.2 Universality Assessment for Diverse Modules
Table 2 demonstrates the core determinants of universality. Although RCAB [11] possesses larger $n$ and $l$ compared to RSTB [13], the elevated $k$ in RSTB deteriorates its transferability. This finding aligns with empirical evidence: RSTB integrates three nested submodules: MSA, MLP, and STL, whereas RCAB employs only the channel attention. Hence, effective migration of RSTB requires the concurrent adaptation of multiple components. Conversely, despite $n_{\mathrm{CRB}}\gg n_{\mathrm{DCRB}}$, DCRB remains more difficult to transfer owing to the inclusion of depth-wise convolution, even though at a macroscopic level $\phi(\text{DCRB})\approx\phi(\text{CRB})$.
@@
Figure 7: Qualitative comparison of remote sensing images for $\times 4$ SISR.
Figure 8: Visualization of the reconstructed results. “Model+” denotes insertion of the original plug-and-play module, “Model++” the UAE-optimized module, and “Model+++” the module further refined by both UAE and CRB. All embedded modules are in accordance with Table 7.
4.3 Comparison Between UAE and Other Metrics
Metrics Evaluation Object Measurement Focus Unit
Parameters Model, Module Storage requirements, Model capacity K($10^{3}$), M($10^{6}$)
FLOPs Model, Module Computational overhead, Inference speed GFLOPs
PSNR, SSIM Model output Reconstruction quality, Structural Similarity dB, Dimensionless ([0,1])
Parameter efficiency Model, Module Performance per parameter Metric per parameter
GPU time Model, Module Inference time on hardware ms
UAE (ours) Module Structural flexibility, Transferability Dimensionless ($0,+\infty]$)
Table 6: Comparison between UAE and other existing metrics.
Figure 9: Spearman Correlation Analysis on UAE. “GPU” represents modules’ inference time on GPU, “PE” stands for parameter efficiency. Fig.LABEL:Spearman_alpha_only, LABEL:Spearman_beta_only, LABEL:Spearman_theta_only present the Spearman correlation analysis between each single component of UAE and other metrics, aiming to investigate which aspects of existing metrics are captured by $k,n,l$ in UAE.
As shown in Table 6, UAE differs from conventional metrics in several aspects. While existing metrics such as PSNR and GPU time primarily focus on performance evaluation at the model/module level, UAE is specifically designed to evaluate the structural flexibility and transferability of individual modules. This enables UAE to capture an essential yet previously unmeasured aspect of module design: its ability to migrate across different networks.
Additionally, given that module transferability is affected by multiple factors, Fig.9 employs the Spearman Correlation Analysis to demonstrate that UAE is a composite metric integrating several metrics. Each single factor $\alpha,\beta,\theta$ exhibits moderate to strong correlation with certain existing metrics. For instance, $\alpha(l)$ correlates highly with CPU time and FLOPs, while $\beta(k)$ has a moderate correlation with parameter efficiency. $\beta$ exhibits a strong negative correlation with PSNR, because modules with very deep nesting levels fail to function independently and must operate within a complete model. However, UAE evaluates the performance of an isolated module to investigate its transferability. As a result, for modules with large $k$, the PSNR performance tends to decrease as the nesting depths increases, leading to a pronounced negative correlation.
The full UAE metric (Fig.LABEL:Spearman_UAE) results demonstrate that UAE is not dominated by a single factor (e.g., 0.72 with parameters, 0.5 with GPU time), but rather offers a more balanced assessment that captures aspects not reflected by existing metrics alone. Therefore, UAE is a novel metric that reveals the combined influence of multiple existing metrics on transferability, an effect not captured by any individual metric.
@@ 4.4 Computational Overhead Analysis
As illustrated in Fig.5, CRB achieves the highest PSNR, while DCRB delivers comparably strong fidelity with the lowest FLOPs, parameter count, and GPU time. Furthermore, CRB outperforms RB in PSNR under equivalent parameters (tested on Potsdam [47]). The convergence profiles in Fig.5 indicate that all universal modules exhibit similar training patterns, with CRB attaining marginally faster convergence within the first 20 epochs, a behavior attributable to the accelerated gradient flow afforded by its additional residual branches.
@@ 4.5 SISR for General Scene Images
Using EDSR [6] as the backbone network, we conduct a combined analysis of module universality and performance, as shown in Table 3. Nearly all of the optimized modules feature better generalizability than baseline RB by slightly sacrificing their universality. Notably, RCAB–despite suffering the greatest universality degradation–fails to achieve the best performance. In contrast, CRB attains the largest generalization gains with minimal impact on universality, marking it as a positively universal module that could be integrated into diverse architectures without additional computational overhead. This finding demonstrates that CRB represents an exceptionally ideal module and indicates that universality and generalizability are not isolated. Although prior work has shown that higher generalization could be attained by sacrificing module transferability via an increased complexity and representational capacity, their precise interplay remains contingent on the specific architecture of plug-and-play modules.
Table 4 summarizes CRB and DCRB performance across multiple SR scales. EDSR augmented by CRB consistently achieves the highest fidelity (PSNR/SSIM), whereas DCRB yields only minor performance degradations—-never exceeding a 5% gap and reaching a minimum of 0.23% under $\times 4$ SR on B100. Remarkably, DCRB comprises just 1.73% of the parameter count of CRB and RB (Table 2), yet under certain conditions delivers approximately 99% of the baseline performance.
As illustrated in Fig.6, we present the $\times 3$ SR results on the B100 test set to visually compare CRB (Ours1) and DCRB (Ours2). Both variants achieve marked clarity enhancements over the LR, yielding similar qualitative reconstructions. The residual images reveal that CRB substantially reduces errors in critical details (e.g., roof eaves in 97033.png). In contrast, DCRB incurs increased residual magnitudes compared to the baseline, particularly within detailed texture areas where errors are more pronounced.
@@ 4.6 SISR For Remote Sensing Images
Table 5 reports the remote-sensing SISR results. Models augmented with CRB consistently outperform their baseline counterparts, demonstrating CRB’s superior universality and its positive impact on model generalizability. In particular, “IPG++CRB” achieves the best performance. In contrast, owing to DCRB’s depthwise separable convolution design, DCRB occasionally underperforms CRB.
@@
4.7 UAE Optimization on Other Basic Modules and Low-Level Tasks
Deblurring Models GOPRO
PSNR SSIM LPIPS
FMANet (CVPR 2024) 27.57 0.8339 0.0783
PPA (ICME 2024) + FMANet 27.71 0.8383 0.0725
$\text{PPA}^{\text{*}}$ + FMANet 27.69 0.8398 0.0719
$\text{PPA}^{\text{**}}$ + FMANet 27.72 0.8387 0.0699
Deraining Models Rain200H
PSNR SSIM LPIPS
NeRD-Rain (CVPR 2024) 30.13 0.9145 0.0430
ELAB (ECCV 2022) + NeRD-Rain 30.19 0.9140 0.0428
$\text{ELAB}^{\text{*}}$ + NeRD-Rain 30.23 0.9143 0.0426
$\text{ELAB}^{\text{**}}$ + NeRD-Rain 30.42 0.9161 0.0422
Denoising Models SIDD-Medium
PSNR SSIM LPIPS
AFM (CVPR 2024) 29.63 0.5843 0.0093
Agent-Attention (ECCV 2024) + AFM 30.61 0.6180 0.0091
$\text{Agent-Attention}^{\text{*}}$ + AFM 31.76 0.6920 0.0089
$\text{Agent-Attention}^{\text{**}}$ + AFM 32.56 0.7490 0.0077
Joint Denoising & Low-Light-Enhancement Models VILNC
PSNR SSIM LPIPS
$\text{ZeroIG}^{{\dagger}}$ (CVPR 2024) 10.17 0.2952 0.4708
ARConv (CVPR 2025) + $\text{ZeroIG}^{{\dagger}}$ 12.13 0.3353 0.3809
$\text{ARConv}^{\text{*}}$ + $\text{ZeroIG}^{{\dagger}}$ 12.23 0.4346 0.3932
$\text{ARConv}^{\text{**}}$ + $\text{ZeroIG}^{{\dagger}}$ 12.50 0.3806 0.3911
Table 7: UAE optimization on different modules and performance evaluation of the optimized models on various low-level tasks. $\text{Module}^{\text{*}}$” represents the module optimized by UAE, and $\text{Module}^{\text{**}}$” represents the CRBs embedded on the basis of UAE optimization. Among the optimized variants, the structural parameters of ZeroIG are moderately reduced to adapt to limited computing resources (marked as $\text{ZeroIG}^{{\dagger}}$), resulting in reduced model performance. However, model performance could still be compared by measuring the relative values of metrics between models.
Firstly, we briefly summarize the models and corresponding UAE optimization methods for the four tasks. Since our UAE optimization stays only within the module, our methods do not affect the inherent plug-and-play characteristics of the modules. We only select appropriate places to embed our optimized components into the framework without changing models’ core architectures.
FMANet (Flow-Guided Dynamic Filtering and Iterative Feature Refinement with Multi-Attention Network) [51] comprises a multi-layer convolutional feature extractor followed by point-wise mappings, emphasizing per-pixel enhancement via stacked convolution blocks. PPA (Parallelized Patch-aware Attention) [52] is a multi-branch convolutional block fused with local/global attention and spatial attention, used for multi-scale context aggregation. During our UAE optimization process, we augment PPA with channel expansion, parallel depth-wise branches and branch fusion followed by spatial attention to increase the representational capacity. CRB is further inserted after $\text{PPA}^{\text{*}}$ output as a post-refinement stage.
NeRD-Rain (Neural Representations for Image Deraining) [53] is a multi-level architecture built from window processing, using local filters and stacked convolutions for deraining. ELAB (Efficient Long-rang Attention Block) [54] is a block combining local filtering and gated multi-window self-attention (GMSA), excelling at local detail refinement. When designing $\text{ELAB}^{\text{*}}$, we add a lightweight local depth-wise branch and SE-style channel reweighting to better use input features, while the GMSA structure is fully preserved. Further, CRBs are applied after $\text{ELAB}^{\text{*}}$ as a bidirectional residual refinement for detail correction.
AFM (Adversarial Frequency Mixup Framework) [55] is a lightweight DnCNN-style backbone, consisting of multiple convolution layers that perform per-pixel noise restoration. Agent-Attention [56] derives adaptive pooled agent tokens and injects agent information into features. We optimize Agent-Attention by increasing heads and agent numbers, and we further embed CRBs inside the agent wrapper, where the attention output is projected and then passed through CRB before normalization and residual additions.
ZeroIG (Zero-Shot Illumination Guided Framework) [57] is a decomposition-based architecture with Enhancer and Denoising modules arranged in a cascaded fashion, which could perform denoising and low-light-enhancement simultaneously. ARConv (Adaptive Rectangular Convolution) [58] is an adaptive resampling convolution module which fixes the drawbacks of convolution operations within a confined square window. We further wrap the ARConv with channel expansion and multiple parallel branches to better leverage the input features. CRBs are embedded in ZeroIG’s high-frequency branch (H3). To be specific, after the denoising step, CRBs are applied for detail refinement.
Table 7 demonstrates that models embedded with the plug-and-play modules exhibit better performance than the baselines, and after our UAE optimization, the performance has generally been further improved. Additionally, models embedded with the modules that are optimized by both UAE and CRB have achieved new SOTAs on nearly all the datasets.
From a qualitative perspective, as shown in Fig. 8, images restored by the optimized models also have better visual quality than the baselines, and most of the details in LQ (Low Quality) images have been recovered. Furthermore, as shown in the deraining part of Fig. 8, although the performance of NeRD-Rain+ is slightly degraded after directly embedding the plug-and-play module into NeRD-Rain, after UAE optimization and CRB enhancement, however, the model performance still surpasses the baseline, providing direct evidence for the effectiveness of our UAE optimization strategy.
Therefore, our optimization strategy is not only applicable to SR tasks, but also other low-level tasks such as denoising, deblurring, and deraining.
4.8 Ablation Studies
4.8.1 Sensitivity Quantification of UAE Parameters
$\alpha(l)$ $\beta(k)$ $\theta(n)$ RB RCAB ConvFFN RSTB GAL DCRB CRB
0.063 0.234 0.094 0.172 0.328 0.125 0.125
0.011 0.028 0.028 0.061 0.061 0.028 0.030
0.045 0.050 0.035 0.046 0.043 0.017 0.045
0.033 0.087 0.062 0.180 0.169 0.030 0.033
0.179 0.743 0.211 0.505 0.902 0.138 0.359
0.046 0.413 0.165 0.675 1.289 0.220 0.091
0.131 1.309 0.372 1.984 3.543 0.244 0.262
Table 8: Variable Ablation of UAE, where $\phi$3 is specified as the UAE form and the module parameters refer to Table 2.
A logarithmic derivative-based approach and elasticity coefficients are proposed to quantify the variable sensitivity. For a general UAE form where $\phi=\Pi_{i}f_{i}(x_{i})$, the logarithmic differentiation goes as Eq.(21)
$d\ln{\phi}=\sum_{i}\frac{\partial\ln{f_{i}}}{\partial\ln{x_{i}}}d\ln{x_{i}}$ (21)
The sensitivity of $x_{i}$ is defined as the absolute elasticity and its normalized form, as shown in Eq.(22). $S_{i}$ represents the percentage change in $\phi$ caused by a 1% change in $x_{i}$. $x_{i}$ and $f_{i}$ respectively stand for the variables and functions in UAE. For instance, $\{<x_{\alpha},f_{\alpha}>\}=\{<l,\alpha(l)>\}$.
$S_{i}=\left|\frac{\partial\ln{\phi_{i}}}{\partial\ln{x_{i}}}\right|=\left|\frac{x_{i}}{f_{i}}\cdot\frac{\partial f_{i}}{\partial x_{i}}\right|,\hat{S_{i}}=\frac{S_{i}}{S_{\alpha}+S_{\beta}+S_{\theta}},i\in\{\alpha,\beta,\theta\}$ (22)
Additionally, we employ the Sharply value to measure the contribution of each variable, which differs from the sensitivity. Sharply value is a concept from cooperative game theory that allocates the total contribution of each participant by averaging their marginal contributions across all combinations. Given that UAE features three variables, the Sharply equations are defined as Eq.(23).
$\begin{cases}C_{\alpha}=\frac{v(\alpha)-v(\varnothing)+v(\alpha,\beta)-v(\beta)+v(\alpha,\theta)-v(\theta)}{3}\\ C_{\beta}=\frac{v(\beta)-v(\varnothing)+v(\alpha,\beta)-v(\alpha)+v(\beta,\theta)-v(\theta)}{3}\\ C_{\theta}=\frac{v(\theta)-v(\varnothing)+v(\theta,\beta)-v(\beta)+v(\alpha,\theta)-v(\alpha)}{3}\end{cases}$ (23)
Figure 10: Variable Ablation of UAE. Fig.LABEL:Variable_Ablation_1 demonstrates the effects of $x,<x,y>,<x,y,z>(x,y,z\in\{k,n,l\})$ on UAE calculation, and Fig.LABEL:Variable_Ablation_2 depicts the variable sensitivity via elasticity coefficients. Fig.LABEL:Variable_Abaltion_3 describes the absolute variable contributions on UAE.
[…]
4.8.2 Nature of CRB Gradient Explosion and Experimental Validation on the $\mathbf{\varepsilon}$ Constraints
$l$ $\varepsilon$ $\varepsilon\leqslant(l/4+\delta_{\text{safe}})$ $\mathcal{L}_{1}^{\text{min}}$ Set5 (PSNR/SSIM) Set14 (PSNR/SSIM) B100 (PSNR/SSIM)
8 1 7.31 27.39/0.8610 24.17/0.7470 23.96/0.7129
2 7.47 27.36/0.8607 24.15/0.7469 23.95/0.7128
3 15.39 21.46/0.5703 20.58/0.5203 21.93/0.5220
12 2 7.57 27.53/0.8643 24.25/0.7503 23.94/0.7126
3 7.50 27.32/0.8601 24.13/0.7468 23.98/0.7143
4 27.06 17.50/0.3236 17.46/0.3062 19.18/0.3427
16 3 8.00 27.41/0.8622 24.18/0.7486 23.95/0.7133
4 7.77 27.46/0.8629 24.21/0.7492 23.96/0.7135
5 101.92 10.53/0.0320 11.22/0.0349 13.00/0.0629
Table 9: Ablation study of $\varepsilon$. Models are trained for 400 epochs on DIV2K dataset, EDSR is used as the backbone. $\delta_{\text{safe}}$ is set to zero during the experiments. $\mathcal{L}_{1}^{\text{min}}$ represents the minimum L1 loss during training.
As shown in Table 9, when $\varepsilon$ exceeds the boundary $l/4+\delta_{\text{safe}}$, module performance significantly deteriorates, and when $\varepsilon\leqslant l/4+\delta_{\text{safe}}$, module performance remains remarkably similar. Therefore, excessive residual connections do not further enhance model performance.
Figure 11: Gradient norm heatmaps of CRB variants during the training process. $l$ represents the module depth and $\varepsilon$ the number of residual connections.
Fig.14 reveals a critical phenomenon validating the constraint $l/4+\delta_{\text{safe}}$. For a cascade of $m=5$ CRBs, when $\varepsilon\leqslant l/4+\delta_{\text{safe}}$ is satisfied, the gradient norms $\|\nabla_{w}L\|$ for the convolution layers within the initial CRB (closest to the input) are smaller than those within the terminal CRB (closest to the output). This is because $J_{l}^{\text{CRB}_{\varepsilon}}$ for each block satisfies $\|J_{l}^{\text{CRB}_{\varepsilon}}\|_{2}>1$ but still remains bounded, preventing gradient vanishing and explosion. The total gradients at the input of block $k$ can be expressed as:
$\|\nabla_{x_{k}}L\|_{2}\propto\|\nabla_{x_{m}}L\|_{2}\cdot\prod_{j=k}^{m-1}\|J_{j}^{\text{CRB}_{\varepsilon}}\|_{2}$ (24)
While the residual structure $\|J_{j}^{\text{CRB}_{\varepsilon}}\|_{2}>1$ mitigates vanishing gradient, the cumulative product over blocks still induces controlled attenuation towards the input, resulting in $\|\nabla_{w_{1}}L\|<\|\nabla_{w_{m}}L\|$, as shown in Fig.LABEL:CRB_8l_2epsilon, LABEL:CRB_12l_3epsilon, LABEL:CRB_16l_4epsilon. This reflects stable, well-behaved back propagation.
When $\varepsilon>l/4+\delta_{\text{safe}}$, however, the gradient norms $\|\nabla_{w}L\|$ for convolution layers within the initial CRB becomes larger than those within the terminal CRB. This is because violating $\varepsilon\leqslant l/4+\delta_{\text{safe}}$ triggers $\|J_{l}^{\text{CRB}_{\varepsilon}}\|>2$. Therefore, gradient at the input of block $k$ scales exponentially as:
$\|\nabla_{x_{k}}L\|_{2}\propto\|\nabla_{x_{m}}L\|_{2}\cdot\prod_{j=k}^{m-1}\|J_{j}^{\text{CRB}_{\varepsilon}}\|_{2}>\|\nabla_{x_{m}}L\|_{2}\cdot 2^{(m-k)}$ (25)
where the exponential amplification dominates the gradient flow. Given that the input block undergoes amplification over Jacobian multiplications, it is obvious that $\|\nabla_{w_{1}}L\|_{2}>\|\nabla_{w_{m}}L\|_{2}$, manifesting the gradient explosion predicted theoretically. Fig.LABEL:explosion1, LABEL:explosion2, LABEL:explosion3 reveal the explosion’s origin in the earliest layers in CRB1, and this abnormal gradient behaviour directly explains the significant performance degradation observed in Table 9.